OpenStudy (anonymous):

A mixture of 40 kg of bird's seed is worth 6 pesos per kg to make up by taking some worth 4.50 pesos per kg and some worth 8.50 per kg. How many kg of each should be taken? Answer please...solution

OpenStudy (anonymous):

4.5x+8.5*y=6*40 x+y=40

OpenStudy (anonymous):

x=25 y=15

OpenStudy (anonymous):

u are welcome

OpenStudy (anonymous):

write a system of equations 4.5x+8.5y=240 x+y=40 I have this system of equations because x kg of the 4.50 bird seed and y kg of 8.50 bird seed makes 240 dollars (40*6) also x kg + y kg equals 40 kg. Solving the system you need 25 kg of the 4.5 peso bird feed and 15 kg of the 8.5 peso bird feed